OpenStudy (anonymous):

need to differentiate twice here, y''. [d^2y/dx^2]

OpenStudy (anonymous):

this is the question:

OpenStudy (anonymous):

Now whats with double differentiation ?is it your assignment

OpenStudy (anonymous):

\[siny - e^x=5x^6\]

OpenStudy (anonymous):

can u pls show the answer as d^2y/dx^2 = aanswer.... nd its kinda my assignment, im just doing practice questions and need some help with double differentiation.

OpenStudy (anonymous):

Now when I differentiate \[cosy \frac{dy}{dx} - e^x = 30x^5\]

OpenStudy (anonymous):

Yea Sure

OpenStudy (anonymous):

Then I double Differentiate \[cosy \frac{d^2y}{dx^2} - siny \frac{dy}{dx} -e^x = 150x^4\]

OpenStudy (anonymous):

maybe solve for \[y'\] and then do it again would be easier. just a thought

OpenStudy (anonymous):

First of all differentiate both sides once, you will get cos y y'-e^x=30x^5 Now differentiate it again

OpenStudy (anonymous):

Tell me what did u get?

OpenStudy (anonymous):

\[\Large{ \frac{d^2y}{dx^2} = \frac{150x^4 + e^x + siny \frac{dy}{dx} }{cosx}}\]

OpenStudy (anonymous):

\[\cos(y)y'-e^x=30x^5\] \[y'=\frac{30x^5+e^x}{\cos(y)}\]

OpenStudy (anonymous):

Thats what I get but wait for satellite

OpenStudy (anonymous):

no your way may be easier. but you have a \[cos(x)\] should be \[\cos(y)\] and also you have a \[\frac{dy}{dx}\] in your answer. but if you solve for it first, then you can solve the whole thing

OpenStudy (anonymous):

Now use Quotient rule to the result shown by Satellite

OpenStudy (anonymous):

i would solve for \[y'\] first, then take the derivative again, and in place of \[y'\] you put \[\frac{30x^5+e^x}{\cos(y)}\]

OpenStudy (anonymous):

give me a second and i will write it out

OpenStudy (anonymous):

ah typo

OpenStudy (anonymous):

an ugly mess is what you get!

OpenStudy (anonymous):

dy/dx should be square in my answer

OpenStudy (anonymous):

and cosy below it not cosx

OpenStudy (anonymous):

\[\Large{\frac{d^2y}{dx^2} = \frac{150x^4 + e^x +siny \times (\frac{dy}{dx})^2}{cosy}}\]

OpenStudy (anonymous):

ok i will type it out, but i am not doing the algebra. first off i get \[y'=\frac{30x^5+e^x}{\cos(y)}\]

OpenStudy (anonymous):

Maybe let satellite confirm

OpenStudy (anonymous):

then i take the derivative of that using quotient rule to get \[\frac{\cos(y)(120x^4+e^x)+(30x^5+e^x)\sin(y)y'}{\cos^2(y)}\]

OpenStudy (anonymous):

that 120 should be 150

OpenStudy (anonymous):

then i put\[y'=\frac{30x^5+e^x}{\cos(y)}\] to get \[\frac{\cos(y)(150x^4+e^x)+(30x^5+e^x)\sin(y)\frac{30x^5+e^x}{\cos(y)}}{\cos^2(y)}\]

OpenStudy (anonymous):

the algebra is on you

OpenStudy (anonymous):

alright, i managed to reach that point as well.... i can do the rest, thanks all of u so much. (was forgetting to substitute the dy/dx into the 2nd part... o.o)

OpenStudy (anonymous):

good luck with algebra!